OpenStudy (anonymous):

|X-10| = 4 -6 or -14 -6 or 14 6 or 14 6 or -14

OpenStudy (anonymous):

I don't understand the question..

OpenStudy (isaidavila):

|x+a|=b -(x+a)=b or (x+a)=b

OpenStudy (kdr1308):

Are you wanting to know the answer to this equation?

OpenStudy (isaidavila):

Use that ro solve the equation

OpenStudy (anonymous):

@Kdr1308 yes what's the solution ?

OpenStudy (kdr1308):

Well, do you know how to solve two step equations?

OpenStudy (anonymous):

Oh nvm I understand it now...this is really simple...I'll just leave this for ^Kdr to explain .-.

OpenStudy (anonymous):

@Kdr1308 give me an example .

OpenStudy (kdr1308):

Okay --> 2x + 3 = 11 -3 -3 2x = 8 ---------- 2 --> x = 4

OpenStudy (anonymous):

Yes i understand that so that's how I'm doing my problem.

OpenStudy (kdr1308):

Yep, but when you put it into 2 separate equations It'll be =4 and =-4 okay?

OpenStudy (anonymous):

So it will be |X-10| = 4 |X-10| = -4 @Kdr1308

OpenStudy (kdr1308):

Yes, correct!

OpenStudy (anonymous):

So the answers are going to be positive or one negative one positive @Kdr1308

OpenStudy (kdr1308):

Not always! Want me to show you?

OpenStudy (anonymous):

Yes please @Kdr1308

OpenStudy (kdr1308):

 Understood?

OpenStudy (anonymous):

Yeah so the answer is 6 or 14

OpenStudy (kdr1308):

Yep, is that all?

OpenStudy (anonymous):

Yes thank you.